OpenStudy (anonymous):

Simply expression with positive exponents: (1/x+h)-(1/x)/h

OpenStudy (campbell_st):

looks like 1st principles is it \[\frac{ \frac{1}{x +h} - \frac{1}{x}}{h} \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Get a common denominator for all terms: 1/x + h -> h/hx + h^2x/hx (1/x)/h is equal to 1/hx Add all terms: h/hx + (h^2)x - 1 all this is over hx (h^2)x + h -1 ------------- hx